A Bibliography of Publications of Brad Bell

Brad Bell University of Washington Applied Physics Laboratory Seattle, Washington, 98105

USA

Tel: +1 206 543 6855 FAX: +1 206 543 6785

E-mail: brad@apl.washington.edu (Internet)

13 October 2017 Version 1.09

Abstract

This bibliography records publications of Brad Bell.

Title word cross-reference

ℓ_{1} [ABBP11].

-Laplace [ABBP11].

1 [Oce83]. 1724120 [Bel10]. 1981 [Oce81]. 1983 [Oce83].

’83 [Oce83].

acoustic [BBO81, BBM83]. age [BF13]. Algorithm [BP91]. Algorithmic [BB08]. applications [Bel10]. Approximating [Bel01]. Approximation [KNB^{+}15]. arrival [BR91]. Asilomar [Che91]. August [Oce83]. Automatic [KNB^{+}15]. baseline [BBO81, BHMS91]. Bayes [PB07, PB07]. Bayesian [PB05, ABBP13b, PBF13]. between [ABBP13b]. blind [PB07]. block [ABBP13a]. book [Bel10]. Boston [Oce81]. Bucy [BBP09]. Burg [BP91].

Calculating [BPW93]. California [Che91]. Carlo [PB08]. chains [PB08]. codes [BBM83, Bel83]. Computers [Che91]. Conference [Che91, Oce81, Che91]. connection [ABBP13b]. connections [ABBP13a]. constrained [BBP09]. convergence [Bel90]. correcting [Bel83]. correction [BBM83]. covariance [PB07].

data [BBM83, BBS94, BBS96]. Deconvolution [PB04, PB05, PB07]. deep [BBO81]. depth [BHMS91]. detection [BBM83]. Differentiation [KNB^{+}15, BB08]. disease [BF13]. Distributed [PBF13]. dynamic [PB08]. eigenfunctions [PB07]. empirical [PB07]. error [BBM83, Bel83]. estimate [Bel01]. Estimating [BP04, BR91, BBS94, BBS96]. estimation [ABBP13b, Bel88, BF13]. evaluation [Bel88].

field [ABBP13b]. fifth [Che91]. filter [BE86, BR91, BC93]. filtering [BHMS91, Bel10]. Francisco [Oce83]. functions [BP04, BB08, BF13, PB05].

gamma [Bel88]. Gauss [BC93, Bel94]. Gaussian [ABBP13b]. Generalized [Bel88]. Global [Bel90, BE86]. GPS [BHMS91, DB94]. Grove [Che91].

Implementation [BBM83]. implicit [BB08]. inequality [BBP09]. infinite [Bel90]. insulin [PB04]. interior [BBP09]. international [Oce81]. inverse [BPW93]. iterated [BC93, Bel94]. iteration [BPW93].

Kalman [ABBP11, ABBP13a, BHMS91, BC93, Bel94, BBP09, Bel10, PBF13].

Laplace [ABBP11, KNB^{+}15]. likelihood [Bel01, BBP09]. linear [PB08]. Long [BBO81, BHMS91]. long-baseline [BHMS91].

marginal [Bel01]. Markov [PB08]. Massachusetts [Oce81]. matched [BE86, BR91]. maximization [BBP09]. MCMC [PB05]. measurement [BP04]. measurements [BHMS91]. Method [DB94, Bel90, BC93, Bel94, BBS94, BBS96]. model [BF13, PB04]. models [Bel01, BP04]. moment [Bel88]. Monte [PB08]. MR [Bel10]. multidimensional [BE86]. multipaths [BE86]. multiple [BBS94, BBS96]. multitapers [BPW93].

network [BR91]. networks [PBF13]. Newton [BC93, Bel94]. non [PB04, PB08]. non-linear [PB08]. non-stationary [PB04]. Nonlinear [BHMS91, BP04, BBP09]. Nonsmooth [Bel84]. November [Che91]. numerical [ABBP13a].

ocean [BBO81, Oce81]. Oceans [Oce81, Oce83]. one [BP04]. Optimal [PB08, BB08]. optimization [Bel84, BE86, Bel90].

Pacific [Che91]. parameter [Bel88]. parameters [BBS94, BBS96, Bel01, BP04]. physical [PB04]. point [BBP09]. Positioning [DB94]. prior [PB07]. Proceedings [Oce83]. programming [Bel84]. pulse [BR91].

quadratic [Bel84].

random [ABBP13b, Bel01]. rate [PB04]. rates [BF13]. real [Bel10]. real-time [Bel10]. record [Che91, Oce81]. relative [BBS94, BBS96]. results [ABBP13a]. review [Bel10]. RKHS [ABBP13b, PB05]. robust [ABBP11].

San [Oce83]. secretion [PB04]. semi [Bel90, PB07]. semi-blind [PB07]. semi-infinite [Bel90]. Separating [BE86]. September [Oce81, Oce83]. sets [BBS94, BBS96]. short [BHMS91]. short-baseline [BHMS91]. Signals [Che91, PB04]. smooth [PB04]. smoother [ABBP11, Bel94, BBP09]. smoothing [ABBP13a, PB08, PBF13]. Software [Bel83]. spectral [BPW93]. stability [ABBP13a]. static [PBF13]. stationary [PB04]. statistical [BF13]. step [BP91]. stochastic [BP04]. successive [Bel84]. system [BBO81]. Systems [Che91, ABBP13a, PB08].

techniques [Bel83, PB05]. telemetry [BBO81, BBM83]. Thomson [BPW93]. time [Bel10, BF13]. times [BR91]. TMB [KNB^{+}15]. tracking [BBO81]. tridiagonal [ABBP13a]. Twenty [Che91]. Twenty-fifth [Che91]. two [BP91].

update [BC93]. using [BP04, PB05, PB07].

values [BB08]. variable [BP04]. variance [PB04]. variances [BBS94, BBS96]. via [PB08].

weighting [BBS94, BBS96]. workplace [Oce81].

References

Aravkin:2011:LRK

[ABBP11] Aleksandr Y. Aravkin, Bradley M.

Bell, James V. Burke, and Gianluigi Pillonetto. An ℓ_{1}-Laplace robust Kalman smoother. IEEE Transactions on Automatic Control, 56(12):2898–2911, 2011. CO-DEN IETAA9. ISSN 0018-9286 (print), 1558-2523 (electronic).

Aravkin:2013:KSB

[ABBP13a] A. Y. Aravkin, B. B. Bell, J. V. Burke, and G. Pillonetto. Kalman smoothing and block tridiagonal systems: new connections and numerical stability results. arXiv.org, March

2013. URL http://arxiv.org/
abs/1303.5237.

Aravkin:2013:CBB

[ABBP13b] A. Y. Aravkin, B. M. Bell, J. V.

Burke, and G. Pillonetto. The connection between Bayesian estimation of a Gaussian random field and RKHS. arXiv.org, Jan-

uary 2013. URL http://arxiv.
org/abs/1301.5288v3.

Bell:2008:ADI

[BB08] Bradley M. Bell and James V. Burke. Algorithmic differentiation of implicit functions and optimal values. In Advances in automatic differentiation, volume 64 of Lect. Notes Comput. Sci. Eng., pages 67–77. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008.

Backes:1983:IED

[BBM83] J. L. Backes, B. M. Bell, and J. B. Miller. Implementation of error detection and correction codes for acoustic data telemetry. In Oceans’83 [Oce83], pages 167– 175. LCCN TC 1505 O331945 1983. Three volumes. IEEE publication number 83CH1972-9.

Backes:1981:LBD

[BBO81] J. L. Backes, B. M. Bell, and L. O. Olson. Long baseline deep ocean acoustic tracking and telemetry system. In Oceans ’81 [Oce81], pages 1–8. LCCN TC1505 .O33193. Two volumes. IEEE publication number 81CH1685-7.

Bell:2009:ICN

[BBP09] Bradley M. Bell, James V. Burke, and Gianluigi Pillonetto. An inequality constrained nonlinear Kalman–Bucy smoother by interior point likelihood maximization. Automatica J. IFAC, 45 (1):25–33, 2009. CODEN AT-CAA9. ISSN 0005-1098 (print), 1873-2836 (electronic).

Bell:1994:RWM

[BBS94] B. M. Bell, J. V. Burke, and A. Schumitzky. A relative weighting method for estimating parameters and variances, in multiple data sets. Computational Statistics & Data Analysis, 1994. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). Submitted.

Bell:1996:RWM

[BBS96] Bradley M. Bell, James V. Burke, and Alan Schumitzky. A relative weighting method for estimating parameters and variances in multiple data sets. Computational Statistics & Data Analysis, 22(2):119–135, 1996. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Bell:1993:IKF

[BC93] Bradley M. Bell and Frederick W. Cathey. The iterated Kalman filter update as a Gauss– Newton method. IEEE Transactions on Automatic Control, 38(2):294–297, 1993. CODEN IETAA9. ISSN 0018-9286 (print), 1558-2523 (electronic).

Bell:1986:SMG

[BE86] B. M. Bell and T. E. Ewart. Separating multipaths by global optimization of a multidimensional matched filter. IEEE Speech and Signal Processing, ASSP-34 (5):1029–1037, 1986.

Bell:1983:STE

[Bel83] B. M. Bell. Software techniques for error correcting codes. Technical Report APL-UW TM-4-83, University of Washington Applied Physics Laboratory, Seattle, WA, USA, 1983.

Bell:1984:NOS

[Bel84] B. M. Bell. Nonsmooth optimization by successive quadratic programming. PhD thesis, University of Washington, 1984.

Bell:1988:GGP

[Bel88] Bradley M. Bell. Generalized gamma parameter estimation and moment evaluation. Communications in Statistics: Theory and Methods, 17(2):507–517, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bell:1990:GCS

[Bel90] Bradley M. Bell. Global convergence of a semi-infinite optimization method. Applied Mathematics and Optimization, 21(1):69– 88, 1990. CODEN AMOMBN. ISSN 0095-4616 (print), 1432-0606 (electronic).

Bell:1994:IKS

[Bel94] Bradley M. Bell. The iterated Kalman smoother as a Gauss– Newton method. SIAM Journal on Optimization, 4(3):626– 636, 1994. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Bell:2001:AML

[Bel01] Bradley M. Bell. Approximating the marginal likelihood estimate for models with random parameters. Applied Mathematics and Computation, 119(1):57– 75, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bell:2010:KFR

[Bel10] Bradley M. Bell. Kalman filtering with real-time applications [book review of MR 1724120]. SIAM Review, 52(2):390–392, 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bell:2013:SME

[BF13] Bradley M. Bell and Abraham D.

Flaxman. A statistical model and estimation of disease rates as functions of age and time. SIAM Journal on Scientific Computing, 35(2):B511–B528, 2013. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Bell:1991:NKF

[BHMS91] B. M. Bell, B. M. Howe, J. A.

Mercer, and R. C. Spindel. Nonlinear Kalman filtering of longbaseline, short-baseline, GPS, and depth measurements. In Chen [Che91], pages 131–136. ISBN 0-8186-2470-1 (paperback), 0-8186-2471-X (microfiche), 0-8186-2472-8 (hardback). LCCN TK 5102.5 A78 1991. IEEE catalog number 91CH31120.

Bell:1991:TSB

[BP91] B. M. Bell and D. B. Percival. A two step Burg Algorithm. IEEE Transactions on Signal Processing, 39(1), 1991.

Bell:2004:EPS

[BP04] Bradley M. Bell and Gianluigi Pillonetto. Estimating parameters and stochastic functions of one variable using nonlinear measurement models. Inverse Problems, 20(3):627–646, 2004. CO-DEN INPEEY. ISSN 0266-5611 (print), 1361-6420 (electronic).

Bell:1993:CTS

[BPW93] B. M. Bell, D. B. Percival, and A. T. Walden. Calculating Thomson’s spectral multitapers by inverse iteration. Journal of Computational and Graphical Statistics, 2(1):119–130, 1993. ISSN 1061-8600 (print), 1537-2715 (electronic).

Bell:1991:MFN

[BR91] B. M. Bell and S. A. Reynolds. A matched filter network for estimating pulse arrival times. IEEE Transactions on Signal Processing, 39(2), 1991.

Chen:1991:CRT

[Che91] Ray R. Chen, editor. Conference record of the Twenty-fifth Asilomar Conference on Signals, Systems and Computers: November 4–6, 1991, Pacific Grove, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-2470-1 (paperback), 0-8186-2471-X (microfiche), 0-8186-2472-8 (hardback). LCCN TK 5102.5 A78 1991. IEEE catalog number 91CH31120.

Dailey:1994:MGP

[DB94] D. J. Dailey and B. M. Bell.

A method for GPS positioning. IEEE Transactions on Aerospace and Electronics, 1994. Submitted.

Kristensen:2015:TAD

[KNB^{+}15] K. Kristensen, A. Nielsen, C. W.

Berg, H. Skaug, and B. Bell. TMB: Automatic differentiation and Laplace approximation. arXiv.org, September 2015.

URL http://arxiv.org/abs/
1509.00660v1.

IEEE:1981:OOI

[Oce81] Oceans 81: The ocean … an international workplace conference record, Boston, Massachusetts, September 16–18, 1981. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. LCCN TC1505 .O33193. Two volumes. IEEE publication number 81CH1685-7.

IEEE:1983:POS

[Oce83] Proceedings, Oceans ’83: San Francisco, August 29–September 1, 1983. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1983. LCCN TC 1505 O331945 1983. Three volumes. IEEE publication number 83CH1972-9.

Pillonetto:2004:DNS

[PB04] Gianluigi Pillonetto and Bradley M. Bell. Deconvolution of nonstationary physical signals: a smooth variance model for insulin secretion rate. Inverse Problems, 20(2):367–383, 2004. CODEN IN-PEEY. ISSN 0266-5611 (print), 1361-6420 (electronic).

Pillonetto:2005:BDF

[PB05] Gianluigi Pillonetto and Bradley M. Bell. Bayesian deconvolution of functions in RKHS using MCMC techniques. In System modeling and optimization, volume 166 of IFIP Int. Fed. Inf. Process., pages 257–268. Kluwer Acad. Publ., Boston, MA, 2005.

Pillonetto:2007:BEB

[PB07] Gianluigi Pillonetto and Bradley M. Bell. Bayes and empirical Bayes semi-blind deconvolution using eigenfunctions of a prior covariance. Automatica J. IFAC, 43(10):1698–1712, 2007. CO-DEN ATCAA9. ISSN 0005-1098 (print), 1873-2836 (electronic).

Pillonetto:2008:OSN

[PB08] Gianluigi Pillonetto and Bradley M. Bell. Optimal smoothing of nonlinear dynamic systems via Monte Carlo Markov chains. Automatica J. IFAC, 44(7):1676–1685, 2008. CODEN ATCAA9. ISSN 0005-1098 (print), 1873-2836 (electronic).

Pillonetto:2013:DKS

[PBF13] Gianluigi Pillonetto, Bradley M. Bell, and Simone Del Favero. Distributed Kalman smoothing in static Bayesian networks. Automatica J. IFAC, 49(4):1001– 1011, 2013. CODEN ATCAA9. ISSN 0005-1098 (print), 1873-2836 (electronic).